OpenStudy (anonymous):

There are several ways to solve system of equations and inequalities, including graphing and using equivalent forms of equations and inequalities within the system. SOLVE USING 2 DIFFERENT METHODS a. 3x-9y=3 b.7x-3y=20 C. y=1/2x-6 6x-3y=-24 5x+3y=16 2x+6y=19

OpenStudy (tkhunny):

Well, you should get to it. Which two methods will you select for each?

OpenStudy (anonymous):

Substitution seems the easiest

OpenStudy (anonymous):

And elimination using addition, subtraction and multiplication

OpenStudy (tkhunny):

Let's see it. No graphing?

OpenStudy (anonymous):

no graphing. I find it too complicated.

OpenStudy (tkhunny):

Okay, for elimination, which equation will you be multiplying by what? I recommend C1 being multiplied by 2 before you do anything else.

OpenStudy (anonymous):

I really need help, because I don't understand how to do this AT ALL! Could you walk me through it step by step? please? I haven't learned this yet. These are my very first questions over this subject.

OpenStudy (anonymous):

Keep in mind that I am only in 7th grade.

OpenStudy (tkhunny):

What do you get when multiplying C1 by 2? Doesn't matter what grade you are in. y=1/2x-6 Multiply everything by 2.

OpenStudy (anonymous):

3 and -12

OpenStudy (anonymous):

y=3x-12 right??

OpenStudy (tkhunny):

2*1 = 2 2* 1/2 = 1 2 * -6 = -12 2y = x - 12 This is also the only one not in standard form. Can you get it to standard form, matching all the others?

OpenStudy (anonymous):

What is standard form?

OpenStudy (tkhunny):

Look at the other five. You have to look at the clues around you. Which variable appears first? Where is the '=' in relation the the variables and constant terms?

OpenStudy (anonymous):

Can you solve one for me? But be sure to show me everything you did to get the answer. Then I can try to do the rest with your help.

OpenStudy (tkhunny):

No. If you do not have sufficient background even to make one equation look like the other five, then no demonstration of a solution will be helpful to you. You must show some ability to solve the problems. So far, you have not done so. Why were you given these problems? Is this a placement exam?

OpenStudy (anonymous):

I was given these problems as in introduction to my system of equations and inequalities unit. If you show me what you did for just ONE... I can do the rest based off of what I saw you do. You would really help me a lot, and help me learn this faster.

OpenStudy (tkhunny):

You need basic algebra skills that you don't seem to have.